Consider the following system of linear equations
x + 3 y = k
4 x + h y = 8
1
Expert's answer
2019-12-31T13:34:11-0500

For a system of equation to have infinite many solutions,

Then,

"\frac{a_1}{a_2}=\frac{b_1}{b_2}=\frac{c_1}{c_2}"

Using the same equation here,

"\frac{1}{4}=\frac{3}{h}=\frac{k}{8}" ;

Hence,"h=12,k=\frac{8}{4}" "=2" .